### 18 Mar 2026 v2

# Training-free Detection of Generated Videos via Spatial-Temporal Likelihoods

Abstract
Following major advances in text and image generation, the video domain has surged, producing highly realistic and controllable sequences. Along with this progress, these models also raise serious concerns about misinformation, making reliable detection of synthetic videos increasingly crucial. Image-based detectors are fundamentally limited because they operate per frame and ignore temporal dynamics, while supervised video detectors generalize poorly to unseen generators, a critical drawback given the rapid emergence of new models. These challenges motivate zero-shot approaches, which avoid synthetic data and instead score content against real-data statistics, enabling training-free, model-agnostic detection. We introduce STALL, a simple, training-free, theoretically justified detector that provides likelihood-based scoring for videos, jointly modeling spatial and temporal evidence within a probabilistic framework. We evaluate STALL on two public benchmarks and introduce ComGenVid, a new benchmark with state-of-the-art generative models. STALL consistently outperforms prior image- and video-based baselines. Code and data are available here.
Abstract
In this supplementary material document, we provide additional implementation details to ensure the full reproducibility of STALL. We also present extended explanations of the statistical tests used to assess the normality of embeddings and the uniformity of the temporal representation features. Furthermore, we include additional experiments and experimental details on all experiments. Next, we provide further details on the newly introduced synthetic dataset, ComGenVid, and important details on the other used benchmarks. We conclude with an efficiency analysis, comparing our method to other zero-shot and supervised methods. The source code, dataset, and pre-computed whitening parameters are publicly available here.
  1. A.
    Reproducibility (Section A).
  2. B.
    Statistical tests (Section B).
  3. C.
    Datasets (Section C).
  4. D.
    Experimental details and additional results (Section D).
  5. E.
    Efficiency analysis (Section E).
1 Introduction
Generative modeling has progressed rapidly across modalities, enabling powerful text and image-generation capabilities built on large language models and diffusion-based image synthesizers [44, 72, 69, 67, 17]. After major breakthroughs in text and image generation, the video domain has undergone a sharp leap forward in the past few years, with highly realistic, controllable video generation models producing long, high-fidelity sequences [23, 38, 77]. These advances unlock strong benefits for creative workflows, content production, and media automation [30, 73]. At the same time, synthetic videos can be misused for misinformation, fraud, impersonation, and intellectual-property violations [4, 52, 43], prompting platforms and regulators to require disclosure of AI-generated content and underscoring the urgency of reliable detection [49, 53]. Unlike deepfake detection, which focuses on manipulation of real content, we address a different problem: detecting fully generated videos, where every frame is synthetic.
In the image domain, early studies mainly relied on supervised classifiers, typically CNN-based models trained to distinguish real from synthetic images using large, labeled datasets [7, 11, 26]. While effective on known generators, these methods require continuous retraining as new generative models emerge and thus generalize poorly to unseen ones [33]. To reduce dependence on synthetic training data, later works explored unsupervised and semi-supervised approaches, leveraging large pretrained models [61, 27]. Recently, zero-shot image detectors have emerged, showing improved robustness and generalization [68, 42, 28, 14]. In this context, zero-shot means no additional training and no generated content available. However, when applied to videos, image detectors assess authenticity only on a per-frame basis. As a result, they ignore temporal dependencies and miss artifacts that emerge across time, such as motion inconsistencies, that are invisible in any single frame.
In the video domain, progress has been more limited. Recent methods predominantly use supervised training to detect generated videos [75, 5, 24, 48, 91], but they inherit the same limitations as supervised image detectors: they require large labeled datasets and generalize poorly to unseen generators. The first zero-shot detector for generated videos is D3 [92], introduced only recently. It analyzes transitions between consecutive frames and relies solely on temporal cues, while ignoring per-frame visual content and spatial information. Moreover, it lacks principled theoretical foundations, relying primarily on empirical hypotheses about real video dynamics. Therefore, a critical gap remains: the need for a mathematically grounded video detector that jointly analyzes spatial content and temporal dynamics.
To address this gap, we introduce STALL, a zero-shot video detector that accounts for both spatial and temporal dimensions when determining whether a video is real or generated (see illustration in LABEL:fig:teaser). Our method leverages a probabilistic image-domain approach [10] and uses DINOv3 [siméoni2025dinov3] to compute image likelihoods. We extend this approach with a temporal likelihood term that captures the consistency of transitions between frames. Unlike prior approaches that are supervised, rely solely on spatial cues (image detectors), or focus exclusively on temporal dynamics [92], our formulation jointly models both aspects and detects inconsistencies that emerge from their interaction (see Figure 2 for qualitative examples). STALL assumes access to a collection of real videos in its pre-processing stage, termed the calibration set. With the abundance of publicly available videos, this is a very mild requirement. The approach is training-free and requires no access to generated samples from any model. The core of the algorithm is based on a new spatio-temporal likelihood model of real videos. This yields a principled measure of how well a video aligns with real-data statistics in space and time.
Our method achieves state-of-the-art performance on two established benchmarks [24, 40] and on our newly introduced dataset comprising videos from recent high-performing generators [62, 36]. We curate this dataset to reflect the newest wave of high-fidelity video models, enabling evaluation on frontier systems. The method is lightweight and efficient, operating without training, and is thus suitable for real-time or large-scale screening pipelines. Across all experiments, it remains robust to common image perturbations, variations in frames-per-second (FPS), and ranges of hyperparameter settings. Our main contributions are as follows:
  • Temporal likelihood. We extend spatial (image-domain) likelihoods to temporal frame-to-frame transitions.
  • Theory-grounded zero-shot video detector. A detector derived from a well-defined theory, which we empirically validate. This provides a principled, measurable tool for analyzing and debugging edge cases.
  • State-of-the-art across benchmarks. We achieve state-of-the-art results on three challenging benchmarks and perform extensive evaluations demonstrating robustness and consistent performance across settings.
  • New benchmark. We release ComGenVid, a curated benchmark featuring recent high-fidelity video generators (e.g., Sora, Veo-3) to support future research.
Figure 2: Qualitative comparison of ZED, D3, and our method (STALL). Each row shows a video clip with natural or unnatural spatial and temporal behavior, together with the corresponding predictions. ZED (spatial-only) misses in cases dominated by temporal inconsistency; D3 (temporal-only) fails when spatial realism is misleading. STALL fuses spatial and temporal likelihoods, yielding robust detection when either modality alone is insufficient. Additional examples with more details are given in Supp. D.7.
2 Background and Related work
Generated image detection. Early work trained supervised CNNs on labeled real and synthetic datasets, sometimes emphasizing hand-crafted artifacts, but generalization to unseen generators was limited [79, 7, 33, 11, 6, 57, 93, 84]. Few-shot and semi or unsupervised variants improved data efficiency by leveraging pretrained features, yet typically retained some dependence on synthetic data or generator assumptions [90, 26, 61, 70, 27]. Zero-shot methods avoid synthetic content exposure by comparing an image to transformed or reconstructed variants [68, 42, 28, 14]. However, these image-only approaches are confined to per-frame spatial cues and ignore cross-frame temporal consistency, leaving them blind to anomalies that only manifest in motion or inter-frame transitions.
Generated video detection. Unlike deepfakes, which edit real footage (e.g., face swaps or lip-sync), we target fully generated content, where the video is synthesized from scratch. Supervised detectors train on labeled real and synthetic videos and report strong in-domain results but struggle in unseen models regime [75, 5]. Recent work also couples new benchmarks with architectures: GenVideo with the DeMamba module [24]; VideoFeedback, which also presents VideoScore (human-aligned automatic scoring) [40]. Parallel efforts explore MLLM-based supervised detectors that provide rationales but still require curated training data and tuning [86, 35]. D3, the first zero-shot video detector, relies on second-order temporal differences, focusing on motion cues [92]. A similar first-order approach is presented in [15]. In contrast, our approach is directly probabilistic and jointly scores spatial (per-frame) and temporal (inter-frame) likelihoods, addressing both appearance and dynamics in a single framework.
Gaussian embeddings and likelihood approximation. Modern visual encoders such as CLIP [66] learn high-dimensional embedding spaces with rich semantic structure. Empirical studies have characterized geometric phenomena in CLIP representations, including the modality gap, narrow-cone concentration [55], and a double-ellipsoid structure [54]. Recent work demonstrates that CLIP embeddings are well-approximated by Gaussian distributions, enabling closed-form image likelihood approximation using whitening without additional training [9]. Whitening has also been shown effective for LLM activations [65]. From a theoretical angle, the Maxwell–Poincaré lemma implies that uniform normalized high-dimensional features have approximately Gaussian projections [31]. This principle has recently been leveraged to show that the InfoNCE objective asymptotically induces Gaussian structure in learned embeddings [8]. Motivated by both empirical evidence and theoretical guarantees, we introduce a normalization step in the temporal embedding space to promote Gaussian statistics and compute faithful likelihood estimates. Additionally, this Gaussian modeling approach extends to other vision encoders [siméoni2025dinov3, 47] and forms the basis of our spatio-temporal video likelihood score.
Figure 3: Method overview. A video is split into frames and encoded into embeddings. The spatial branch scores the likelihood of each frame embedding; the temporal branch normalizes inter-frame differences and scores their likelihood analogously. The two scores are then fused into a unified measure that separates AI-generated from real videos.
3 Preliminaries
We now introduce the mathematical tools and notations used throughout the paper. These concepts form the basis of our likelihood formulation and will be applied in the method section (Section 4).
3.1 Whitening transform and Gaussian likelihood
Notation. Let $\mathcal{X}=\{x_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$ and let $\mathbf{X}\in\mathbb{R}^{d\times N}$ be the column-stacked matrix. Define the sample mean $\mu=\tfrac{1}{N}\sum_{i=1}^{N}x_{i}$ and centered vectors $\hat{x}_{i}=x_{i}-\mu$, with $\mathbf{\hat{X}}=[\hat{x}_{1},\dots,\hat{x}_{N}]$. The empirical covariance is $\mathbf{\Sigma}=\tfrac{1}{N}\mathbf{\hat{X}}\mathbf{\hat{X}}^{\top}$.
Whitening transform. We seek a linear transform $W\in\mathbb{R}^{d\times d}$ that admits:
$W^{\top}W\;=\;\mathbf{\Sigma}^{-1}.$ (1)
The whitening matrix is not unique: if $W$ satisfies Equation 1, then so does $RW$ for any orthogonal $R$. A common choice is PCA-whitening. Let the eigen-decomposition be $\mathbf{\Sigma}=V\Lambda V^{\top}$ with eigenvectors $V$ and eigenvalues $\Lambda=\mathrm{diag}(\lambda_{1},\dots,\lambda_{d})$. The PCA-whitening matrix is
$W_{\text{PCA}}\;=\;\Lambda^{-\tfrac{1}{2}}V^{\top}.$ (2)
Given a vector $x$, the whitened representation is $y\;=\;W(x-\mu)$ and the whitened data matrix is
$\mathbf{Y}\;=\;W\,\mathbf{\hat{X}}$. (3)
Whitened embeddings have zero mean and identity covariance.
Likelihood approximation. Under the zero-mean and identity-covariance properties, if the whitened coordinates follow a Gaussian distribution, then $y\sim\mathcal{N}(0,I_{d})$. Given this isotropic Gaussian model, the log-likelihood is:
$\ell(y)\;=\;\log p(y)\;=\;-\tfrac{1}{2}\Big(d\log(2\pi)+\|y\|_{2}^{2}\Big),$ (4)
where $\|y\|_{2}^{2}=y^{\top}y$. Given an embedding $x$, the whitened norm $\|W(x-\mu)\|_{2}^{2}$ thus provides a closed-form likelihood proxy when the Gaussian assumption holds.
3.2 Asymptotic Gaussian projections
When vectors are uniformly distributed on the unit sphere in $\mathbb{R}^{d}$, their coordinates behave approximately Gaussian. The Maxwell-Poincaré lemma [31, 32] formalizes this: if $u\sim\mathrm{Unif}(\mathbb{S}^{d-1})$, then for each coordinate,
$\sqrt{d}\,u_{j}\xrightarrow[d\to\infty]{}\mathcal{N}(0,1).$ (5)
More generally, for high-dimensional vectors with nearly uniform directions and concentrated norms, any fixed low-dimensional linear projection is well-approximated by a Gaussian. Supplementary Material (Supp.) Section B.3 details the lemma and convergence rates.
Algorithm 1 STALL (Generated video detector)
1:Encoder $E$, calibration set $\mathcal{C}$, test video $v=\{f_{t}\}_{t=1}^{T}$
2:
3:Calibration (pre-processing)
4:From $\mathcal{C}$: extract all frames per video, encode $\{x_{i,t}\}$.
5:Compute spatial statistics $(\mu,W)$Equation 2), using a single frame per video.
6:Compute the temporal statistics $(\mu_{\Delta},W_{\Delta})$Equation 2) using all normalized inter-frame differences for each video.
7:Record calibration distributions of $s_{C_{\mathrm{sp}}}$ and $s_{C_{\mathrm{temp}}}$.
8:
1:Inference (test-time)
2:Compute $x_{t}=E(f_{t})$; for frames $\{f_{t}\}_{t=1}^{T}$ of $v$.
[…]
7:return $s_{\mathrm{video}}\leftarrow\tfrac{1}{2}\big(\mathrm{perc}(s_{\mathrm{sp}})\;+\;\mathrm{perc}(s_{\mathrm{temp}})\big)$.
4 Method: STALL
We propose STALL (Spatial-Temporal Aggregated Log-Likelihoods), a zero-shot detector that jointly scores videos via a spatial likelihood over per-frame embeddings and a temporal likelihood over inter-frame transitions. A high-level overview of the method is shown in Figure 3, and Algorithm 1 summarizes the procedure. Detailed algorithms for all method steps are provided in Supp. Section A.1.
Notation. Let $\mathcal{C}=\{c^{(i)}\}_{i=1}^{N}$ denote a collection of videos. A video $c\in\mathcal{C}$ consists of $T$ frames, written as $c=\{f_{t}\}_{t=1}^{T}$. Each frame $f_{t}$ is mapped to a $d$-dimensional embedding using a vision encoder $E$, yielding $x_{t}=E(f_{t})\in\mathbb{R}^{d}$.
4.1 Spatial likelihood
Prior work [10] in the image domain observed that whitened CLIP embeddings are well-approximated by standard Gaussian coordinates, as verified on MSCOCO [56], using Anderson–Darling (AD) and D’Agostino–Pearson (DP) normality tests [2, 29]. Therefore, the norm in the whitened space correlates with the likelihood of an image. We extend this result to the video setting by extracting frame-level embeddings from real video datasets. We apply the whitening procedure discussed above (Section 3.1), and assess Gaussianity with the same tests, evaluating multiple encoders. Under this Gaussian assumption, per-frame spatial likelihoods follow the closed-form log-likelihood in Equation 4. Details and results are in Supp. Section B.1.
We estimate spatial likelihood statistics using a calibration set of $n$ real videos (see Section 4.4). This step involves no training and is computed a priori only once. It consists of estimating real-data statistics, which remain fixed throughout inference. At inference time, for a test video $v=\{f_{t}\}_{t=1}^{T}$, each frame $f_{t}$ is encoded as $x_{t}=E(f_{t})$, whitened to $y_{t}$ using Equation 3, and assigned a log-likelihood $\ell(y_{t})$ according to Equation 4.
4.2 Temporal likelihood
Spatial likelihoods score frames independently; they do not assess how transitions evolve across time. To capture motion consistency, we examine the embedding space and model frame-to-frame differences, $\Delta_{t}\;=\;x_{t+1}-x_{t}$.
Normalization Induces Gaussianity. Empirically, the raw transition vectors $\Delta_{t}$} are not well modeled by a Gaussian distribution (see Supp. Section B.1). We observe that these high-dimensional transitions exhibit two key properties: (1) Variable magnitudes; their norms vary substantially across samples; and (2) Random directions; their orientations are approximately spanned in a uniform manner, since the underlying video motions are arbitrary and thus lack any preferred direction; see Supp. Section B.1 for empirical validation. In high-dimensional spaces, uniformly distributed directions on the sphere behave similarly to Gaussian samples when projected onto any axis, as established by the Maxwell–Poincaré lemma [31, 32] (Section 3.2). To obtain a stable probabilistic model, we normalize each transition vector as $\tilde{\Delta}_{t}=\frac{\Delta_{t}}{\lVert\Delta_{t}\rVert}$, placing all transition directions on the unit sphere. Empirically, these normalized transitions exhibit Gaussian-like behavior, see illustration in Figure 5 and quantitative results in Supp. Section B.1.
(a) Pearson correlation
(b) Spearman correlation
Figure 4: Correlations among spatial and temporal aggregation methods. Values computed on VATEX [81], which is not used in our evaluations. When all individual likelihood detectors perform reasonably well on the evaluated benchmarks (see Supp. Section D.2), lower correlations are desirable.
Corner case: if two consecutive frames are identical ($x_{t+1}=x_{t}$), their transition vector is $\Delta_{t}=x_{t+1}-x_{t}=0$. Such transitions carry no temporal information and are deterministically discarded from the temporal likelihood computation. If all frames in a video are identical, i.e., $f_{1}=f_{2}=\dots=f_{T}$, the input effectively degenerates to a single image. In this case no temporal score is defined and the detector falls back to the spatial likelihood $s_{\mathrm{spatial}}(V)$, which analyzes the image domain.
Using the calibration set of real videos, we collect all normalized transition vectors $\{\tilde{\Delta}_{t}\}$ and compute their empirical mean $\mu_{\Delta}$ and covariance $\Sigma_{\Delta}$. At inference time, in a manner analogous to the spatial likelihood, we whiten the normalized transitions, $z_{t}$, using Equation 3, and compute their log-likelihoods $\ell(z_{t})$ according to Equation 4. This yields the temporal log-likelihood of each transition in the video. Generated videos often exhibit unnatural motion, resulting in transitions with low likelihood under this model.
Table 1: Zero-shot detection results. Comparison of image- and video-based detectors on three benchmarks. Best in each row is bold; second best is underlined. Our method achieves the highest average performance on all benchmarks and leads on most generators individually. It is also the only method that detects consistently across all models, maintaining AUC >0.5>0.5.
Benchmark Model Image Detectors Video Detectors AEROBLADE [68] RIGID [42] ZED [28] D3 (L2) [92] D3 (cos) [92] STALL (Ours) AUC AP AUC AP AUC AP AUC AP AUC AP AUC AP VideoFeedback [40] AnimateDiff [37] $0.57$ $0.55$ 0.73 0.74 $0.65$ $0.62$ $0.49$ $0.49$ $0.61$ $0.57$ 0.83 0.86 Fast-SVD [12] $0.52$ $0.51$ $0.54$ $0.56$ $0.45$ $0.48$ $0.76$ $0.77$ 0.80 0.79 0.89 0.89 LVDM [41] 0.88 0.90 $0.65$ $0.57$ $0.76$ $0.70$ $0.42$ $0.49$ $0.31$ $0.41$ 0.86 0.89 LaVie $0.50$ $0.50$ 0.71 0.73 $0.29$ $0.37$ $0.51$ $0.47$ $0.49$ $0.46$ 0.81 0.83 ModelScope [78] $0.60$ $0.56$ $0.66$ 0.62 0.69 $0.59$ $0.51$ $0.52$ $0.42$ $0.46$ 0.81 0.83 Pika [64] $0.44$ $0.46$ $0.54$ $0.54$ $0.39$ $0.47$ 0.83 0.84 0.81 0.81 $0.78$ $0.80$ Sora [62] $0.65$ $0.62$ $0.43$ $0.44$ $0.56$ 0.62 $0.62$ $0.56$ 0.67 $0.58$ 0.81 0.82 Text2Video[51] $0.67$ $0.63$ 0.70 0.68 $0.55$ $0.49$ $0.15$ $0.33$ $0.22$ $0.36$ 0.83 0.83 VideoCrafter2 [23] $0.60$ $0.58$ 0.80 $0.76$ $0.53$ $0.50$ $0.69$ $0.71$ $0.80$ 0.79 0.93 0.94 ZeroScope [20] 0.78 0.78 $0.65$ $0.59$ $0.70$ $0.62$ $0.35$ $0.45$ $0.35$ $0.44$ 0.78 0.81 Hotshot-XL [46] $0.20$ $0.34$ $0.51$ $0.58$ $0.44$ $0.45$ 0.64 0.67 $0.60$ $0.62$ 0.79 0.80 Average $0.58$ $0.58$ 0.63 0.62 $0.54$ $0.54$ $0.54$ $0.57$ $0.55$ $0.57$ 0.83 0.85 GenVideo [24] Crafter [22] $0.64$ $0.65$ $0.71$ $0.66$ $0.55$ $0.56$ 0.79 0.82 $0.76$ $0.79$ 0.82 0.80 Gen2 [34] $0.56$ $0.59$ $0.70$ $0.67$ $0.51$ $0.58$ 0.88 0.90 0.88 0.90 0.88 $0.89$ Lavie [82] $0.58$ $0.59$ 0.77 0.76 $0.39$ $0.42$ $0.68$ $0.68$ $0.67$ $0.68$ 0.85 0.84 ModelScope [78] $0.60$ $0.60$ $0.62$ $0.59$ $0.61$ $0.57$ 0.63 0.64 $0.60$ $0.63$ 0.78 0.78 MorphStudio [59] 0.74 0.73 0.74 $0.69$ $0.60$ $0.60$ $0.66$ $0.71$ $0.64$ $0.69$ 0.83 0.84 Show_1 [89] $0.48$ $0.50$ $0.53$ $0.52$ $0.45$ $0.47$ 0.76 0.80 $0.75$ $0.79$ 0.82 0.80 Sora [62] $0.73$ $0.70$ $0.49$ $0.48$ $0.71$ 0.79 0.75 $0.75$ $0.74$ $0.74$ 0.79 0.80 WildScrape [85] $0.49$ $0.53$ $0.61$ $0.59$ $0.55$ $0.57$ 0.65 0.69 $0.64$ 0.69 0.72 $0.68$ HotShot-XL [46] $0.31$ $0.39$ 0.64 0.65 $0.47$ $0.46$ $0.56$ $0.64$ $0.54$ $0.62$ 0.79 0.78 MoonValley [58] $0.75$ $0.78$ $0.72$ $0.66$ $0.63$ $0.72$ 0.81 0.82 0.81 0.82 $0.72$ $0.75$ Average $0.59$ $0.61$ $0.65$ $0.63$ $0.55$ $0.57$ 0.72 0.74 $0.70$ $0.74$ 0.80 0.80 ComGenVid (ours) Sora [62] 0.72 0.67 $0.53$ $0.55$ $0.58$ $0.59$ $0.68$ $0.65$ $0.68$ $0.65$ 0.84 0.85 VEO3 [36] $0.67$ $0.62$ $0.62$ $0.63$ $0.52$ $0.55$ 0.79 $0.76$ 0.79 0.78 0.86 0.87 Average $0.69$ $0.64$ $0.57$ $0.59$ $0.55$ $0.57$ 0.73 0.71 0.73 0.71 0.85 0.86 All Benchmarks Average $0.62$ $0.61$ $0.61$ $0.59$ $0.57$ $0.58$ 0.64 0.65 0.64 0.65 0.82 0.82
Figure 5: Temporal embedding coordinates. Raw coordinates of temporal embeddings (frame differences) are not Gaussian; after normalization, each coordinate is approximately Gaussian (full histogram comparison in Supp. B.2.1).
4.3 Unified score
We compute likelihood scores for each frame (spatial) and each frame-to-frame transition (temporal). We first aggregate each list separately and then combine the two aggregates into a single video-level score. We evaluate standard aggregation operators: minimum, maximum, and mean, on a set of real videos and measure the cross-domain correlations induced by each choice (Figure 4). Combining the minimum of one domain with the maximum of the other yields the lowest correlation, indicating complementary information. Accordingly, we use the minimal temporal likelihood and the maximal spatial likelihood per video. The method is robust to this selection; detection results for all combinations are reported in Supp. Section D.2.
Percentile scoring. Because spatial and temporal likelihoods lie on different scales, we avoid raw magnitudes and compare each score relative to real data, so decisions reflect how typical a video is under the calibration distribution. We set aside the spatial and temporal scores from the calibration set and, at inference, convert a test score $s$ into a rank-based percentile by counting how many calibration scores $s_{1},\ldots,s_{n}$ satisfy $s_{i}\leq s$ and dividing by $n$: $\mathrm{perc}(s)\;=\;\frac{1}{n}\,\bigl|\{\,i:s_{i}\leq s\,\}\bigr|.$ We compute these percentiles separately for the spatial and temporal scores.
Unified video score. The final video score aggregates the percentile-normalized components:
$s_{\mathrm{video}}(v)\;=\;\frac{1}{2}\Big(\text{perc}_{\mathrm{sp}}(v)+\text{perc}_{\mathrm{temp}}(v)\Big).$ (6)
Percentile normalization makes both terms scale-free and less sensitive to extreme OOD values. In  Section 5.3, we ablate each component (spatial/temporal) alone and cross-component fusion (average vs. product) and find robustness across choices. Each component is individually discriminative, and the unified score performs best.
4.4 Calibration set
We use a calibration set of real videos to compute whitening statistics and percentile ranges, aligning with zero-shot detection: no generated samples are used at any point, and “in-distribution” is defined solely by real data. The calibration set is disjoint from all evaluation benchmarks and any other data used elsewhere in this paper, ensuring no overlap or leakage. This is not a limitation: every detector must define a decision boundary, and real-only calibration provides a principled, data-driven anchor for both spatial and temporal likelihoods. Ablations are provided in Sec. 5.3.
(a) Latency comparison.
(b) Performance comparison.
(c) Calibration set ablation.
Figure 6: Comparison of detectors and calibration data. We compare our method against six detectors: three image-based [68, 42, 28] and three video-based [92, 1, 5], for performance and efficiency. (a) Inference latency per video. (b) Average AUC across all three benchmarks; our method is both high-performing and efficient. (c) GenVideo [24] results using different datasets as the calibration set, showing that same-distribution calibration is only slightly better, indicating robustness to the calibration choice (see Supp. D.6.6).
5 Evaluations
5.1 Experimental settings
Datasets. We evaluate our detector on two benchmarks spanning real and generated videos. VideoFeedback [40] contains $\sim$33k generated videos from 11 text-to-video models [64, 51, 23, 78, 82, 37, 41, 46, 20, 12, 16] and $\sim$4k real videos drawn from two datasets [25, 3]. GenVideo [24] (test set) comprises $\sim$8.5k generated videos from 10 generative sets [78, 59, 58, 45, 89, 34, 22, 82, 16] and $\sim$10k real videos from a single dataset [87]. Across both benchmarks, the generative models constitute a diverse collection of diffusion-based text-to-video systems. Additionally, we present ComGenVid, a set of $\sim$3.5k generated videos from recent commercial models Veo3 and Sora [36, 62], designed to stress cross-model generalization. We pair these with $\sim$1.7k real videos sampled from [21]. For all evaluations, we subsample to use equal numbers of real and generated videos (determined by the smaller class in each split) to ensure fair metric comparisons. A complete breakdown of generative models, video counts, and dataset composition is given in Supp. Section C.
(a) Calibration set size ablation.
(b) Image perturbations.
Figure 7: Robustness to calibration set size and image perturbations. (a) Varying calibration-set size between 1k to 34k; each size is resampled 5 times and we report mean AUC $\pm$ standard deviation. (b) Robustness to four common image perturbations applied randomly to frames at five severity levels; our method maintains high separation across perturbation type and intensity. Both experiments are performed on GenVideo [24].
Metrics. We report Area Under the ROC Curve (AUC) and Average Precision (AP). AUC measures the ability of the detector to separate real and generated videos by integrating the ROC curve (true-positive-rate vs. false-positive-rate across thresholds), while AP summarizes the precision–recall trade-off for the positive (generated) class.
Implementation details. We use available official implementations for baselines: AEROBLADE [68] and D3 (both L2 and cosine similarity variants, see Supp. Section A.4), and the supervised detectors T2VE [1] and AIGVdet [5] (official weights and code). For RIGID [42] and ZED [28], we reimplemented the authors’ methods following the paper’s specifications (see Supp. Section A.2). Image detectors operate per-frame, and we report the mean score over frames. In all experiments we encode frames using DINOv3 [siméoni2025dinov3] for our method, and use a fixed calibration set that is built from 33k real videos from VATEX [81]. This dataset is completely separate from any data used for evaluations. We conduct ablations on calibration set size and dataset, encoder model and method components in the next section.
Data curation and evaluation protocol. Following standard protocols [5, 92], we standardize inputs to 8 or 16 frames. For fair comparison, we sample all evaluated videos at 8 FPS and 2 s duration (16 frames). The only exceptions are HotShot-XL and MoonValley [46, 58], which generate 1 s videos; for these we compare against real 1 s videos at 8 FPS (8 frames). Image detectors operate per frame and the average score over all frames is evaluated. We report results under this default setting (Supp. A.3) and provide an ablation on FPS/duration sensitivity in the next section.
(a) Step size ablation.
(b) Length of video ablation.
(c) FPS ablation.
Figure 8: Temporal ablations. Our method remains robust under all temporal variations: (a) temporal step size for likelihood computation; (b) video length; (c) frame rate (FPS).
5.2 Results
Benchmark evaluations. Table 1 reports zero-shot results across all three benchmarks. Our method achieves the highest average performance on each benchmark and attains the best per-generator results in most cases; when not the top method, it remains competitive. Notably, all other methods produce AUC values below 0.5 for some generators, indicating an inverted decision boundary; detectors that fit one model misclassify many examples from others. Our method does not exhibit this failure mode and maintains consistent separation between real and generated samples. In Figure 6(b) we also include supervised video detectors; our zero-shot method still outperforms them, even though they are partially trained on the evaluated generators.
Efficiency. We measure each method’s inference time, measured per video (16 frames input), results are reported in Figure 6(a). STALL, together with D3 [92] and RIGID [42] are the fastest methods with 0.49s, 0.5s and 0.6s respectively. ZED [28] and T2VE [1] show double latency (0.92s, 0.97s) while AEROBLADE [68] and AIGVdet [5] demonstrate increased latency. Our method is relatively lightweight (Supp. Section E) making it highly efficient.
5.3 Ablation study
Calibration set. We study the effect of calibration set size and dataset. We examine different datasets as the source of the calibration set in  Figure 6(c). We compare VATEX [81] with a combination of Kinetics-400 and PE datasets [50, 13], samples from the real data of VideoFeedback and samples from the real data of Genvideo (different samples than the evaluation set). We add a combination of all four options. Using data from the same distribution as tested (GenVideo) results in only a slightly higher performance. Other options remain competitive, demonstrating robustness to dataset selection. We then vary the calibration set size (using VATEX [81]) between 1k and 34k and report mean AUC and standard deviation over 5 sampling iterations. Results are in  Figure 7(a) showing that only at very small sizes (less than 5k) results drop significantly. For whitening, we use one frame per video and all frame-to-frame transitions. We test using a single transition per video and find that results remain identical (Supp. Section D.3).
Embedder comparison. We evaluate our detector with multiple vision encoders: DINOv3 [siméoni2025dinov3], lightweight MobileNet [47], and ResNet-18 [39], as well as video encoders VideoMAE [74] and ViCLIP [83]. Video encoders produce a single embedding per video, and we compute likelihood directly on this vector. Table 2 shows that image encoders perform strongly, even with older, lightweight backbones like MobileNet and ResNet-18. Video encoders, by contrast, perform poorly. Collapsing an entire video to a single embedding discards frame-wise and transition statistics, undermining both spatial and temporal likelihood modeling.
Table 2: Backbone encoder ablation.
Image encoders Video encoders DINOv3 [siméoni2025dinov3] MobileNet-v3 [47] ResNet-18 [39] ViCLIP-L/14 [83] VideoMAE [74] AUC 0.81 0.82 0.79 0.59 0.61
Robustness analysis. We test robustness by applying standard image corruptions to video frames: JPEG compression, Gaussian blur, resized crop, and additive noise, at five severity levels. Perturbations are applied only at inference while keeping the calibration set unchanged. Results in Figure 7(b) show strong robustness across perturbation types up to the highest intensity levels; implementation details and examples in Supp. Section D.5. We also evaluate robustness to temporal perturbations in Supp. Section D.4.
We further vary the input FPS and the temporal step size between frames ($\Delta_{t}$) to assess sensitivity to motion sparsity and sampling rate, and also test different video durations. Results in Figure 8 show that our method remains robust across all temporal settings. Evaluations settings of these experiments are in Supp. Section D.6.
Finally, we assess performance using higher-order temporal differences: while temporal transitions capture first-order changes, higher orders model more complex motion dynamics. As shown in Supp. Section D.1, all orders exhibit high correlation and yield nearly identical results.
Component analysis. We assess three variants: (i) spatial-only, (ii) temporal-only, and (iii) the full model combining both. For each, we report results with raw likelihoods and with percentile-ranked scores. We also test standard aggregations (min, max, mean). Results show that either single-domain detector performs well, the combined detector performs best, and performance is robust to the choice of aggregation (see Supp. Section D.2).
6 Conclusion
We introduce STALL, a zero-shot detector for fully generated videos that fuses spatial (per-frame) and temporal (inter-frame) likelihoods in a single probabilistic framework. Our method is training-free, uses no generated samples, and relies solely on real videos to define reference distributions for both spatial and temporal statistics. Across multiple benchmarks, including recent frontier models such as Sora and Veo3, our approach consistently outperforms prior supervised and zero-shot image/video detectors. It is also efficient and robust to spatial and temporal perturbations, calibration data size and source, and aggregation choices. As this field continues to develop rapidly, there remains room for improvement; nevertheless, our results highlight that modeling the statistical structure of real videos is a promising path for robust detection.
Acknowledgments
We would like to acknowledge support by the Israel Science Foundation (Grant 1472/23) and by the Ministry of Innovation, Science and Technology (Grant 8801/25).
A Reproducibility
A.1 Detailed algorithms
To ensure complete reproducibility, we provide full implementation details, including detailed algorithms for whitening (Algorithm 2), scores (Algorithm 3), (Algorithm 4), calibration (Algorithm 5), and inference (Algorithm 6).
A.1.1 Notation
  • $\mathcal{C}=\{c^{(i)}\}_{i=1}^{N_{c}}$: Calibration set of $N_{c}$ real videos
  • $c^{(i)}=\{f_{t}^{(i)}\}_{t=1}^{T}$: Calibration video $i$ consisting of $T$ frames.
  • $v=\{f_{t}\}_{t=1}^{T}$: query video with $T$ frames.
  • $E:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{d}$: Image vision encoder.
  • $x_{t}=E(f_{t})\in\mathbb{R}^{d}$: Embedding of frame $f_{t}$.
  • $\Delta_{t}=x_{t+1}-x_{t}$: Temporal difference between consecutive embeddings.
  • $\tilde{\Delta}_{t}=\Delta_{t}/\|\Delta_{t}\|_{2}$: $\ell_{2}$-normalized temporal difference.
  • $\mu,W$: Mean and whitening matrix for spatial embeddings
  • $\mu_{\Delta},W_{\Delta}$: Mean and whitening matrix for temporal embeddings
  • $s_{\text{spat}},s_{\text{temp}}$ spatial and temporal scores.
A.1.2 Algorithms
Algorithm 2 Compute Whitening Transform
1:Embeddings $\mathcal{X}=\{x_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$
[…]
6:Eigendecomposition: $\mathbf{\Sigma}=V\Lambda V^{\top}$
[…]
8:return $\mu,W$
Algorithm 3 Compute Spatial Score
1:Frame embeddings $\{x_{t}\}_{t=1}^{T}$, parameters $(\mu,W)$
[…]
5:return $s_{\text{spat}}$
Algorithm 4 Compute Temporal Score
1:Frame embeddings $\{x_{t}\}_{t=1}^{T}$, parameters $(\mu_{\Delta},W_{\Delta})$
[…]
7:return $s_{\text{temp}}$
Algorithm 5 STALL Calibration
1:Calibration set $\mathcal{C}=\{c^{(i)}\}_{i=1}^{N_{c}}$ of $N_{c}$ videos, each with $T$ frames, encoder $E$
2:Encode all frames from calibration set:
3:for $i=1$ to $N_{c}$ do
4:  $x_{t}^{(i)}\leftarrow E(f_{t}^{(i)})$ for $t=1,\ldots,T_{i}$
5:end for
6:Compute spatial whitening parameters:
[…]
8:for $i=1$ to $N_{c}$ do
9:  Sample one frame: $x\sim\text{Uniform}(\{x_{t}^{(i)}\}_{t=1}^{T_{i}})$
10:  $\mathcal{X}_{\text{spat}}\leftarrow\mathcal{X}_{\text{spat}}\cup\{x\}$
11:end for
[…]
13:Compute temporal whitening parameters:
[…]
15:for $i=1$ to $N_{c}$ do
16:  for $t=1$ to $T_{i}-1$ do
17:   $\Delta_{t}\leftarrow x_{t+1}^{(i)}-x_{t}^{(i)}$
18:   $\tilde{\Delta}_{t}\leftarrow\frac{\Delta_{t}}{\|\Delta_{t}\|_{2}}$
19:   $\mathcal{X}_{\text{temp}}\leftarrow\mathcal{X}_{\text{temp}}\cup\{\tilde{\Delta}_{t}\}$
20:  end for
21:end for
[…]
23:Compute calibration score distributions:
24:for $i=1$ to $N_{c}$ do
25:  $s_{\text{spat}}^{(i)}\leftarrow\textsc{Compute Spatial Score}(\{x_{t}^{(i)}\}_{t=1}^{T_{i}},\mu,W)$ $\triangleright$ Algorithm 3
26:  $s_{\text{temp}}^{(i)}\leftarrow\textsc{Compute Temporal Score}(\{x_{t}^{(i)}\}_{t=1}^{T_{i}},\mu_{\Delta},W_{\Delta})$ $\triangleright$ Algorithm 4
27:end for
[…]
30:return $\mu,W,\mu_{\Delta},W_{\Delta},\mathcal{S}_{\text{spat}},\mathcal{S}_{\text{temp}}$
Algorithm 6 STALL Inference
1:Test video $v=\{f_{t}\}_{t=1}^{T}$, encoder $E$, calibration parameters $\mu,W,\mu_{\Delta},W_{\Delta},\mathcal{S}_{\text{spat}},\mathcal{S}_{\text{temp}}$
2:// Encode all frames from test video
[…]
4:Compute spatial and temporal scores:
[…]
7:Compute percentile ranks from calibration distributions:
[…]
10:Combine percentiles for final detection score:
[…]
12:return $s_{\text{video}}$
A.2 Implementation details
All of our experiments are conducted using Intel Core i9-7940X CPU and NVIDIA GeForce RTX 3090 GPU. For our method we use DINOv3 [siméoni2025dinov3] as our encoder, available at DINOv3 repository. For all competing methods, we rely on the official implementations when available, otherwise, we implement the corresponding baselines.
For AEROBLADE [68], we use the official implementation available at AEROBLADE repository. Since no official implementation is available for RIGID [42], we implement the method using the DINO [19] model. As there is no official implementation for ZED [28] either, we follow the lossless compression setup based on SReC [18]. ZED introduces four separate criteria, and we report results for the $\Delta^{01}$ criterion. For AIGVDet [5] and T2VE [1], we use the official code and pretrained weights released by the authors, available at AIGVDet repository and T2VE repository, respectively. For D3 [92], we rely on the official implementation at D3 repository, and use DINOv3 as the encoder.
A.3 Comparison Methodology
A.3.1 Video Preprocessing
In all of our evaluations, we filter out videos that are shorter than 2 seconds or have a frame rate below 8 FPS to ensure sufficient temporal coverage and quality. After filtering, we subsample frames to achieve a target frame rate of 8 FPS. Given an original frame rate $f_{\text{orig}}$ and target rate $f_{\text{target}}=8$ FPS, we compute the sampling ratio $r=f_{\text{orig}}/f_{\text{target}}$ and select every $r$-th frame (approximately).
Specifically, we maintain a continuous position that advances by $r$ at each step, selecting the frame at the rounded position:
$i_{0}=0,\quad i_{j}=\text{round}(r\cdot j),\quad j=1,2,\ldots$
subject to $i_{j}<N$, where $N$ is the total number of frames of the original video.
When $f_{\text{orig}}$ is perfectly divisible by $f_{\text{target}}$ (i.e., $r$ is an integer), this reduces to uniform sampling of every $r$-th frame. For non-integer ratios, this approach selects frames with approximately uniform temporal spacing that best approximates the target frame rate. The corresponding Python implementation is provided below:
def downsample_frames(num_frames, current_fps, target_fps=8):
"""Downsample frame indices to achieve target fps."""
ratio = current_fps / target_fps
indices = []
j = 0
while True:
frame_idx = round(ratio * j)
if frame_idx >= num_frames:
break
indices.append(frame_idx)
j += 1
return indices
Unless stated otherwise, all videos in our experiments were sampled at 8 FPS and truncated to 2 seconds, yielding 16 frames per video.
A.3.2 Pairwise Comparison Protocol
We conduct systematic pairwise comparisons between synthetic videos from each generative model and authentic videos. To ensure fair metric comparisons and to address data imbalance, we implement a balanced sampling procedure where we use equal numbers of real and generated videos, determined by the smaller class in each split. For each generative model $M_{i}$ with $N_{i}$ synthetic videos, we sample exactly $N_{i}$ authentic videos from our real video datasets. Specifically, in the Videofeedback benchmark [40], the real videos are drawn from two datasets [25, 3]. To prevent bias from over-representation of any specific real dataset source, we sample an equal number of videos from both sources such that their total equals $N_{i}$.
A.4 D3 Baseline Evaluation
We identified two systematic differences between D3’s [92] official evaluation protocol and ours that explain the performance gap reported in the main paper.
A.4.1 Unbalanced test set
When fewer than 1000 synthetic videos are available, 1000 real videos are compared against fewer synthetic samples (on GenVideo [24], this is most notably the case with Sora [62], which contains only 56 samples). Average Precision (AP) is sensitive to class imbalance and tends to inflate. Our pairwise comparison protocol (Section A.3.2) uses equal numbers of real and generated videos to ensure unbiased evaluation.
A.4.2 FPS upsampling by frame duplication
In the GenVideo  [24] benchmark, the MSR-VTT [87] real videos provided on ModelScope are stored at only 3 FPS. D3’s official code brings these to 8 FPS by duplicating frames. This duplication applies exclusively to real videos, because the generated videos in GenVideo are already at 8 FPS or higher. The resulting temporal redundancy inflates detection scores for methods that rely on inter-frame differences. In our evaluation, we download high-frame-rate MSR-VTT videos and uniformly downsample them to 8 FPS, which removes this artifact. for more details, see  Sections C.2 and A.3.1.
A.4.3 D3 Ablation Results
In our main experiments, we evaluated D3 using the same embedder as our method (DinoV3 [siméoni2025dinov3]) for consistency. For completeness, we now report D3 results using X-CLIP16 [60] as embedder, as conducted in the original D3 implementation, which yields a small performance increase. Table 3 shows how D3 mean AP on GenVideo varies across evaluation settings: choice of embedder, test-set balancing and sampling. Each protocol difference individually inflates AP, and their combination produces a large gap relative to our evaluation.
Table 3: D3 [92] mean AP across evaluation settings on GenVideo. The two protocol differences (class imbalance and FPS duplication) each inflate AP; together they fully explain the gap between D3’s reported numbers and our evaluation.
Real video Sampling Downsample Upsample (Duplication) FPS $\sim 27\rightarrow 8$ $3\rightarrow 8$ Embedder Balanced Unbalanced Balanced Unbalanced X-CLIP16 [60] 0.78 0.85 0.97 0.98 DinoV3 [siméoni2025dinov3] 0.74 0.83 0.94 0.96
B Normality measures and tests
B.1 Normality and uniformity tests
To assess whether the embeddings are approximately Gaussian, we apply two classical normality tests, Anderson–Darlin  [2] and D’Agostino–Pearson [29], performed on each coordinate of the embeddings independently. Each test measures the normality of the one-dimensional vector input. Below we elaborate on each test and present results.
B.1.1 Anderson-Darling normality test
The Anderson-Darling (AD) test [2] can be viewed as a refinement of the classical Kolmogorov-Smirnov (KS) goodness-of-fit test [71], designed to put more weight on discrepancies in the tails of the distribution. Given a sample $X=\{x_{1},x_{2},\dots,x_{n}\}$ and a target cumulative distribution function (CDF) of $F$ (in our case, a normal CDF with parameters estimated from the data), the AD statistic is defined as
$A^{2}=-n-\sum_{i=1}^{n}\frac{2i-1}{n}\bigl[\ln F(x_{i})+\ln\bigl(1-F(x_{n+1-i})\bigr)\bigr],$ (7)
where $x_{1}\leq\dots\leq x_{n}$ are the ordered sample values, $F(x)$ is the CDF of the reference normal distribution, and $n$ is the sample size. Larger values of $A^{2}$ indicate stronger deviations from normality. these values are compared to tabulated critical values to decide whether to reject the Gaussian assumption.In our setting, we follow the conventional threshold $A^{2}<0.752$ as evidence to accept normality. To preform this test, we used stats.anderson function from scipy python package.
B.1.2 D’Agostino-Pearson Test
The D’Agostino-Pearson (DP) test [29] evaluates departures from normality by combining information about sample skewness and kurtosis. Let $X=\{x_{1},x_{2},\dots,x_{n}\}$ be a univariate sample and let $\mu=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ denote its mean. Let $m_{i}=\frac{1}{n}\sum_{j=1}^{n}(x_{j}-\mu)^{i}$ be the $i$-th centeral moment. The skewness $g_{1}$ and kurtosis $g_{2}$ are defined as:
$g_{1}$ $=\frac{m_{3}}{m_{2}^{3/2}}=\frac{\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{3}}{\left[\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{2}\right]^{3/2}}$ (8)
$g_{2}$ $=\frac{m_{4}}{m_{2}^{2}}-3=\frac{\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{4}}{\left[\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{2}\right]^{2}}-3$ (9)
These two statistics are then transformed into approximately standard normal variables $Z_{1}$ and $Z_{2}$, and the DP test statistic is
$K^{2}=Z_{1}^{2}+Z_{2}^{2}.$ (10)
Under the null hypothesis of normality, $K^{2}$ approximately follows a chi square distribution with two degrees of freedom, so the $p$-value is
$p=1-F_{\chi^{2}_{2}}(K^{2}),$ (11)
where $F_{\chi^{2}_{2}}$ is the cumulative distribution function of $\chi^{2}_{2}$. Positive $g_{1}$ indicates right-skewed data and negative $g_{1}$ left-skewed data, while positive $g_{2}$ corresponds to heavy tails and negative $g_{2}$ to light tails. As $g_{1}$ and $g_{2}$ approach zero, the test statistic $K^{2}$ decreases and the $p$-value increases, which is consistent with normality; in practice we treat $p>0.05$ as compatible with a Gaussian distribution. We performed this test using the stats.normaltest function from the scipy Python package.
B.1.3 Results
To obtain stable estimates of normality, we randomly sample $40$ independent groups of $250$ embeddings each from the VATEX [81] calibration set, for both frame embeddings and frame embedding differences. As described and used in the main paper, we employ DINOv3 [siméoni2025dinov3] as the frame level embedder, whose embedding space is $1024$ dimensional. For every group and every coordinate, we apply the Anderson Darling (AD) and D’Agostino Pearson (DP) normality tests. We then aggregate the outcomes in two separate ways: (i) we average the test statistics across all coordinates and groups, and (ii) we compute the fraction of coordinates whose average statistic satisfies the normality thresholds ($A^{2}<0.752$ for AD, $p>0.05$ for DP). As summarized in Table 4, raw frame embeddings $x_{t}$ show high proportions of coordinates passing both tests, and whitening further increases these fractions. In contrast, raw temporal differences $\Delta_{t}$ are strongly non-Gaussian, with essentially no coordinates passing either criterion. After $\ell_{2}$ normalization, temporal differences $\tilde{\Delta}_{t}$ exhibit high pass rates, and an additional whitening step $z_{t}$ yields almost all coordinates satisfying both normality thresholds.
Table 4: Normality tests. Results of Anderson Darling (AD) and D’Agostino Pearson (DP) normality tests for different representations on the VATEX [81] calibration set. “Avg Score” is the mean test statistic across embedding coordinates and groups, “Threshold” specifies the acceptance condition for normality, and “Normal Features” is the percentage of coordinates satisfying that condition.
Representation Test Avg Score Threshold Normal Features $(\uparrow)$ approximately gaussian?
Raw embeddings $x_{t}$ AD 0.4750 <0.752<0.752 96.3%
DP 0.3931 >0.05>0.05 99.2%
Whitened embeddings $y_{t}$ AD 0.5034 <0.752<0.752 98.2%
DP 0.3207 >0.05>0.05 99.3%
Transition vector $\Delta_{t}=x_{t+1}-x_{t}$ AD 3.3992 <0.752<0.752 0.0%
DP 0.0093 >0.05>0.05 0.0%
$\ell_{2}$ Normalized transition vector $\tilde{\Delta}_{t}$ AD 0.4134 <0.752<0.752 98.4%
DP 0.4752 >0.05>0.05 99.9%
Whitened $\ell_{2}$ normalized transition vector $z_{t}$ AD 0.4119 <0.752<0.752 99.6%
DP 0.4648 >0.05>0.05 100.0%
B.2 Histogram comparisons
B.2.1 Raw vs. normalized temporal differences
To qualitatively illustrate these findings, Figure 9 presents histograms of temporal differences for the first four embedding dimensions of DINOv3 [siméoni2025dinov3], computed from all adjacent frame pairs in the VATEX [81] calibration set. The left column shows raw temporal differences (Raw $\Delta_{t}$), which exhibit significant deviations from the overlaid Gaussian distributions. In contrast, the right column shows $\ell_{2}$-normalized temporal differences (Normalized $\Delta_{t}$), which align closely with their corresponding Gaussian fits. This visual demonstration confirms the quantitative results in Table 4: $\ell_{2}$ normalization transforms non-Gaussian temporal differences into approximately Gaussian distributions.
Figure 9: Raw vs. Normalized Temporal Difference Histograms. Histogram comparison of raw temporal differences (Raw $\Delta_{t}$, left) versus normalized temporal differences (Normalized $\Delta_{t}$, right) for the first four dimensions of DINOv3 [siméoni2025dinov3] embeddings, computed from all adjacent frame pairs in the VATEX [81] calibration set. Red curves show Gaussian distributions fitted using the empirical mean and variance of the data (i.e., a moment-matched Gaussian). Even under this moment-matched fit, raw differences exhibit clear deviations from Gaussianity, while normalized differences closely match Gaussian distributions across all dimensions.
B.2.2 Real and AI-generated videos embeddings
To further illustrate the statistical properties of the embedding space, we visualize the univariate feature distributions of embeddings extracted from real and generated videos. Figure 10 shows histograms for the first four dimensions of DINOv3 [siméoni2025dinov3] embeddings, computed from randomly sampled frames from real and generated videos in the GenVideo [24] dataset.
For each dimension, we overlay a Gaussian distribution fitted using the empirical mean and variance of the data (i.e., a moment-matched Gaussian). The empirical distributions closely follow the corresponding Gaussian curves for both real and generated samples. While the means and variances differ slightly between real and generated content, the overall shapes remain approximately Gaussian. These visualizations provide additional qualitative support for modeling the embedding dimensions using Gaussian statistics, which underlies the likelihood formulation used in our method.
Figure 10: Real and AI-generated videos embeddings. Histograms of selected embedding dimensions from DINOv3 [siméoni2025dinov3] features extracted from randomly sampled frames in the GenVideo [24] dataset. The left column corresponds to real videos, and the right column corresponds to generated videos. Red curves show Gaussian distributions fitted using the empirical mean and variance of each feature dimension.
B.3 Maxwell Poincare Lemma
Lemma 1 (Maxwell-Poincaré [31]).
Let $U_{d}$ be uniform on $\mathbb{S}^{d-1}$ and fix $k\in\mathbb{N}$. Then
$\sqrt{d}(U_{d,1},\ldots,U_{d,k})\;\Rightarrow\;\mathcal{N}(0,I_{k})\qquad(d\to\infty).$ (12)
The rate at which this convergence occurs was quantified by Diaconis and Freedman [32]:
Theorem 2 (Maxwell-Poincaré convergence rate[32]).
If $1\leq k\leq d-4$, then
$d_{\mathrm{TV}}\!\big(\sqrt{d}(U_{d,1},\ldots,U_{d,k}),Z\big)\;\leq\;\frac{2(k+3)}{d-k-3}\;,$ (13)
where $Z\sim\mathcal{N}(0,I_{k})$.
Note that both Lemma 1 and Theorem 2 extend naturally to arbitrary coordinate selections, or more generally to any $k$-dimensional orthonormal projection of $U_{d}$. When norms exhibit concentration behavior, an analogous result can be established:
Lemma 3 (Maxwell–Poincaré with norm concentration).
Let $U_{d}$ be as in Lemma 1 and fix $k\in\mathbb{N}$. Let $z_{d}=r_{d}U_{d}$ with $r_{d}\geq 0$. If $r_{d}\xrightarrow[d\to\infty]{\mathsf{P}}r_{0}$, then
$r_{d}\,\sqrt{d}\,(U_{d,1},\ldots,U_{d,k})\;\Rightarrow\;\mathcal{N}\!\big(0,\,r_{0}^{2}I_{k}\big)\qquad(d\to\infty).$ (14)
This is obtained by combining the Maxwell–Poincaré limit with radial concentration and applying Slutsky’s theorem [76].
Figure 11 visualizes the distribution of the first coordinate of $U_{d}$, a random vector uniformly distributed on $\mathbb{S}^{d-1}$, for several dimensions $d$. As $d$ increases, this distribution approaches the standard normal distribution.
Figure 11: Uniformity on the sphere and Gaussian distribution. Histograms of the first coordinate of $U_{d}$, where $U_{d}$ is uniformly distributed on the $d$-dimensional unit sphere $\mathbb{S}^{d-1}$, shown for several values of $d$ and overlaid with the $\mathcal{N}(0,1)$ density.
Figure 12 shows histograms of cosine similarities over all unordered pairs of $3k0$ randomly sampled normalized temporal differences $\tilde{\Delta}_{t}\subset\mathbb{R}^{1024}$ from the VATEX [81] calibration set and of $3k$ points drawn uniformly from $\mathbb{S}^{1023}$.
Figure 12: Uniformity on the sphere. Cosine similarity distributions over all unordered pairs of $3k$ randomly sampled normalized temporal differences $\tilde{\Delta}_{t}$ from the VATEX [81] calibration set and $3k$ points drawn uniformly from the unit sphere.
C Datasets
C.1 ComGenVid benchmark
for full statistics of our ComGenVid benchmark, refer to table 5
MSVD Dataset.
We obtained the complete MSVD dataset [21] from the sarthakjain004 Kaggle repository using the kaggle command line interface:
kaggle datasets download -d sarthakjain004/msvd-clips --unzip
We sampled 1.7k videos with 2 seconds length from this dataset.
The complete set of sampled videos is listed in msvd_sampled_videos.csv.
Generative Model Videos.
To ensure fair comparison, we sampled 1.7k videos from both Sora [16] and Veo3 [36] to match the MSVD dataset size.
Veo3 Sampling. We randomly selected 1.7k videos from the ShareVeo3 dataset [80], available on the WenhaoWang Hugging Face repository, and can be download with the following python script:
# pip install huggingface_hub[hf_xet]
from huggingface_hub import hf_hub_download
for i in range(1, 51):
hf_hub_download(
repo_id="WenhaoWang/ShareVeo3",
[…]
repo_type="dataset"
The complete list of sampled videos is provided in veo3_sampled_videos.csv.
Sora Sampling. We manually collected 1.7k videos from distinct users on the OpenAI Sora public explore feed. The complete list of sampled videos is provided in sora_sampled_videos.csv.
Table 5: Overview of video sources and characteristics in the ComGenVid dataset.
Video Source Type Length Range Length (Mean±Std) Resolution Number of Pixels (Mean±Std) FPS (Mean±Std) Total Count MSVD [21] Real 2-60s 9.68±6.27s 160×112-1920×1080 0.29±0.35M 29.1±8.6 1700 Sora [62] Fake 4-20s 6.01±2.26s 480×480-720×1080 0.36±0.05M 30.0±0.0 1700 VEO3 [36] Fake 8s 8.00±0.00s 1280×720 0.92±0.00M 24.0±0.0 1700 Total Count - - - - - - 5100
C.2 GenVideo
We obtained the AI-generated videos for GenVideo [24] from the official ModelScope collection. Because the MSR-VTT [87] real videos available on ModelScope are limited to 3 FPS, we downloaded the original MSR-VTT dataset from the khoahunhtngng Kaggle repository to access higher-frame-rate versions.
All evaluated videos are uniformly sampled at 8 FPS for a 2-second duration (16 frames). The only exceptions are HotShot-XL and MoonValley [46, 58], which produce clips shorter than 2 seconds, these are compared against real 1-second videos at 8 FPS (8 frames). Generative models that produce less than 2-second clips are excluded from the ablation study. Because MSR-VTT contains far more real videos (10k) than any generative model, we selected 1.4k videos from it, following our pairwise comparison protocol (Section A.3.2). Comprehensive GenVideo benchmark statistics are provided in Table 6.
Table 6: Comprehensive statistics of the GenVideo [24] dataset used in our evaluations.
Video Source Type Length Range Length (Mean±Std) Resolution Number of Pixels (Mean±Std) FPS Range FPS (Mean±Std) Total Count MSR-VTT [87] Real 10-30s 14.80±5.04s 320×240 0.08±0.00M 10-30 27.3±3.3 1400 Crafter [22] Fake 2s 2.00±0.00s 1024×576 0.59±0.00M 8 8.0±0.0 188 Gen2 [34] Fake 4s 4.00±0.00s 896×504-1408×768 0.77±0.31M 24 24.0±0.0 1380 Lavie [82] Fake 2-3s 2.27±0.27s 512×320 0.16±0.00M 8-24 16.0±8.0 1400 ModelScope [78] Fake 4s 4.00±0.00s 448×256-1280×720 0.70±0.36M 8 8.0±0.0 700 MorphStudio [59] Fake 2s 2.00±0.00s 1024×576 0.59±0.00M 8 8.0±0.0 700 Show_1 [89] Fake 4s 3.62±0.00s 576×320 0.18±0.00M 8 8.0±0.0 700 Sora [62] Fake 9-60s 16.78±10.96s 512×512-1920×1088 1.43±0.65M 30 30.0±0.0 56 WildScrape [85] Fake 2-251s 7.41±18.79s 256×256-2286×1120 0.48±0.41M 8-45 16.9±9.7 529 HotShot  [46] Fake 1s 1.00±0.00s 672×384 0.26±0.00M 8 8.0±0.0 700 MoonValley [58] Fake 1.82s 1.82±0.00s 1184×672 0.80±0.00M 50 50.0±0.0 626 Total Count - - - - - - - 8379
C.3 VideoFeedback
We gather the Videofeedback  [40] dataset from the official Hugging Face repository. We evaluate only videos that are at least 2 seconds long, except for HotShot-XL [46], which generates 1 s clips and is therefore compared against real 1 s videos at 8 FPS (8 frames). in the original paper  [40], each clip is assigned a dynamic-degree score (1–4) indicating how clearly its motion can be distinguished from a static image. We retain only the highest-scoring videos (levels 3–4). for the full statistics of the VideoFeedback benchmark, refer to Table 7.
Table 7: Comprehensive statistics for the VideoFeedback [40] benchmark used in our evaluations.
Video Source Type Length Range Length (Mean±Std) Resolution Number of Pixels (Mean±Std) FPS Range FPS (Mean±Std) Total Count DiDeMo [3] Real 3s 3.00±0.00s 352×288-640×1138 0.27±0.09M 8 8.0±0.0 1861 Panda70M [25] Real 2-3s 2.38±0.48s 384×288-640×360 0.23±0.01M 8 8.0±0.0 1861 AnimateDiff [37] Fake 2s 2.00±0.00s 512×512 0.26±0.00M 8 8.0±0.0 992 Fast-SVD [12] Fake 3s 3.00±0.00s 768×432 0.33±0.00M 8 8.0±0.0 959 LVDM [41] Fake 2s 2.00±0.00s 256×256 0.07±0.00M 8 8.0±0.0 2973 LaVie [82] Fake 2s 2.00±0.00s 256×160-512×320 0.13±0.06M 8 8.0±0.0 2789 ModelScope [78] Fake 2s 2.00±0.00s 256×256 0.07±0.00M 8 8.0±0.0 3722 Pika [64] Fake 3s 3.00±0.00s 768×640 0.49±0.00M 8 8.0±0.0 1906 Sora [62] Fake 2-3s 2.73±0.27s 512×512-1920×1088 1.25±0.58M 8 8.0±0.0 898 Text2Video [51] Fake 2s 2.00±0.00s 256×256 0.07±0.00M 8 8.0±0.0 3722 VideoCrafter2 [23] Fake 2s 2.00±0.00s 512×320 0.16±0.00M 8 8.0±0.0 3543 ZeroScope [20] Fake 3s 3.00±0.00s 256×256 0.07±0.00M 8 8.0±0.0 2022 Hotshot-XL  [46] Fake 1s 1.00±0.00s 512×512 0.26±0.00M 8 8.0±0.0 2736 Total Count - - - - - - - 29984
C.4 VATEX (Calibration set)
We obtained the VATEX dataset [81] from the khaledatef1’s Kaggle repositories. The dataset is distributed across three parts: Vatex 1, Vatex 2, and Vatex 3. We downloaded all parts using the Kaggle command-line interface:
kaggle datasets download -d khaledatef1/vatex0110 --unzip
kaggle datasets download -d khaledatef1/vatex01101 --unzip
kaggle datasets download -d khaledatef1/vatex011011 --unzip
For the complete statistics of VATEX [81] calibaration set, see Table 8.
Table 8: Comprehensive statistics of the VATEX [81] calibration set.
Video Source Type Length Range Length (Mean±Std) Resolution Number of Pixels (Mean±Std) FPS Range FPS (Mean±Std) Total Count VATEX [81] Real 2-10s 9.68±1.10s 128×88-720×1280 0.44±0.37M 8-30 27.2±5.0 33976 Total Count - - - - - - - 33976
C.5 Other datasets
We collected approximately 1.5k real videos from Kinetics400 [50] and PE [13]. The Kinetics400 clips were taken from the test split available in the cvdfoundation repository, while the PE videos were downloaded from the Facebook PE Hugging Face page. These datasets were used exclusively for the calibration set ablation experiment reported in Fig. 6(c) in the main paper. A detailed specification of this calibration set configuration is provided in Table 9. A complete list of the sampled videos is provided in pe_kinetics400_sampled_videos.csv.
Table 9: Comprehensive statistics of the Kinetics+PE calibration set.
Video Source Type Length Range Length (Mean±Std) Resolution Number of Pixels (Mean±Std) FPS Range FPS (Mean±Std) Total Count Kinetics400 [50] Real 2-10s 9.62±1.19s 128×96-1280×720 0.55±0.38M 24-30 26.9±3.0 1496 PE [13] Real 5-60s 16.49±9.37s 608×254-608×1152 0.21±0.03M 24-60 37.7±14.5 1500 Total Count - - - - - - - 2996
D Additional results and experimental details
Unless stated otherwise, all ablation studies are conducted on the GenVideo [24] benchmark, restricted to generative-model videos sampled at 8 FPS and 2 seconds duration (16 frames). The only exceptions are HotShot-XL and MoonValley [46, 58], which generate videos shorter than 2 seconds and are therefore omitted from our ablation experiments. We use DINOv3 [siméoni2025dinov3] as the frame-level embedder and the VATEX [81] dataset as the calibration set.
D.1 Derivative order ablations
We isolate the effect of higher-order temporal differences by keeping the entire STALL pipeline fixed and changing only the temporal derivative order $D\in\{1,2,3,4\}$. For each video $v=\{f_{t}\}_{t=1}^{T}$, we extract frame-wise embeddings $E(v)\in\mathbb{R}^{T\times d}$, compute the $D$-th order finite-difference trajectory along time, and apply frame-wise $\ell_{2}$ normalization. Importantly, we fit a separate whitening transform for each temporal derivative order, yielding parameters $(\mu_{\Delta=i},W_{\Delta=i})$ for every $i\in D$, estimated on the corresponding derivative trajectories from the VATEX [81] calibration set. The temporal log-likelihood sequence for each video is aggregated using the same statistic as in the main STALL score, and we then average this temporal percentile with the spatial log-likelihood percentile to obtain a single score for each derivative order. An example implementation of the derivative and normalization computation is shown below:
def differences_vec(features):
return features[:, 1:, :] - features[:, :-1, :]
def temporal_diff_with_order(features, derivative_order: int):
Apply a temporal finite-difference operator of the given order
to frame-wise embeddings and L2-normalize the result.
Args:
frame_embeddings: array of shape [N, T, d]
derivative_order: positive integer order of the derivative
Returns:
Array of shape [N, T - derivative_order, d].
if derivative_order < 1:
raise ValueError("derivative_order must be positive")
for _ in range(derivative_order):
features = differences_vec(features)
norms = np.linalg.norm(features, axis=-1, keepdims=True) + 1e-8
return features / norms
Ablation results across all derivative orders, including AUC, Pearson correlation, and Spearman correlation, are summarized in Figure 13. All derivative orders are strongly correlated and produce very similar performance.
(a) AUC - different derivative orders.
(b) Pearson correlation.
(c) Spearman correlation.
Figure 13: Ablations on temporal derivative order. STALL performance for different temporal derivative orders $D\in\{1,2,3,4\}$: (a) AUC across orders, (b) Pearson correlation between scores from different orders, and (c) Spearman correlation between scores from different orders.
D.2 spatial and temporal aggregation methods
D.2.1 Components ablation
We analyze the contribution of the spatial and temporal branches by evaluating three detector variants: a spatial-only model, a temporal-only model, and a combined model that fuses both branches. For each variant, we consider raw scores and percentile-ranked scores, and for the combined model we also test mean-based and product-based fusions of the spatial and temporal percentile scores. For every configuration, we report the average AP and AUC across all three benchmarks. See results in Fig. 14.
(a) AP - different components
(b) AUC - different components
Figure 14: Components ablation. Bar plots of spatial-only, temporal-only, and combined detectors, using both raw scores and percentile-ranked scores. Spatial and Temporal denote single-branch detectors, and Combined refers to the fused Spatial and Temporal detector. Configurations labeled with “P” use percentile-ranked scores, where “P” stands for percentile. “Avg” stands for average of the two scores and “Prod” stands for the product of the two scores.
D.2.2 Aggregation ablation
To analyze the effect of the frame-level aggregation operators, we fix our pipeline as defined in the main paper and vary only the frame-level aggregation used within each branch. We sweep over min, mean, and max for both the spatial and temporal components. For each spatial and temporal aggregation pair, we compute the average AUC and AP across all three benchmarks and report the resulting scores as heatmaps in Supp. Fig. 15.
(a) AP for different spatial and temporal aggregation pairs.
(b) AUC for different spatial and temporal aggregation pairs.
Figure 15: Frame-level aggregation ablation. Heatmaps of AP and AUC, for different combinations of spatial and temporal frame-level aggregation operators. Rows correspond to the spatial aggregation and columns to the temporal aggregation, while all other parts of the pipeline follow the default configuration described in the main paper.
D.3 Sampling a single frame difference for temporal whitening
In this ablation, we test whether temporal whitening requires all frame-to-frame transitions or can be reliably estimated from a single transition per video. We keep the spatial branch (and its whitening transform) fixed, and re-compute the temporal whitening statistics $(\mu_{\Delta},W_{\Delta})$ twice on the calibration set: (i) using all normalized frame differences from all videos, and (ii) using only one randomly sampled frame difference per video. Re-evaluating the combined score under these two temporal-whitening calibration variants yields essentially the same performance, with average AUC 0.8110 and average AP 0.8046 for (i) and average AUC 0.8105 and average AP 0.8044 for (ii) (Pearson correlation 0.9994, Spearman correlation 0.9992 between (i) and (ii)), indicating that our temporal whitening is robust to the number of transitions used from each video for its estimation.
D.4 Temporal likelihood under temporal perturbations
We analyze the temporal likelihood under realistic perturbations: reversing frame order (rewind), shuffling consecutive frames, and inserting black or white frames, representing data transmission issues [88]. As shown in Figure 16, the method is robust to reversal and shuffling, since the score depends only on statistics of adjacent-frame differences, which are largely preserved under these operations. In contrast, inserting a flash frame introduces an abrupt temporal inconsistency that strongly reduces the likelihood. The experiment is conducted on 400 real videos from MSR-VTT [87]. The temporal likelihood is stable under common temporal distortions but responds strongly to abrupt temporal anomalies.
Figure 16: Temporal scores under temporal perturbations. Each condition applies a different perturbation to 400 real MSR-VTT videos: Original (no perturbation), Reversed (frames in reverse order), Shuffle (consecutive frames shuffled), Black flash (one black frame inserted mid-video), and White flash (one white frame inserted mid-video). The temporal likelihood is largely unaffected by reversal and shuffling, since adjacent-frame difference statistics are preserved; however, abrupt flash frames cause a strong drop in the score, with white flashes producing a larger degradation than black flashes.
D.5 Image perturbation experiment details
To explore robustness to perturbations, we apply four standard image corruptions to GenVideo [24] frames at inference time only, keeping the calibration set fixed. We use a reduced GenVideo subset with 250 videos per generative model and corrupt every video with each perturbation type at all five predefined severity levels. We then measure the impact of each corruption setting on detection performance. We additionally include level 0, which corresponds to uncorrupted frames (no perturbation). As shown in Fig. 7(b) of the main paper, STALL maintains strong separation across all perturbation types and severities.
The specific parameter settings for each perturbation type and level are summarized in Tab. 10. Gaussian blur severity is controlled by the blur radius $r$, with larger $r$ producing stronger smoothing. JPEG compression reduces the image quality parameter, where lower quality introduces stronger compression artifacts. The random resized crop perturbation is parameterized by scale ranges, where a random crop is sampled at each application: sometimes removing more content and sometimes less. Higher severity levels use wider ranges with smaller minimum scales, increasing the chance of more aggressive crops. Gaussian noise severity is adjusted by increasing the standard deviation of zero-mean noise added to each frame, which progressively degrades fine texture while preserving global structure.
Table 10: Image perturbation details. These perturbations are used in the robustness ablation (Sec. D.5). Levels 1–5 span the full range of corruption strength for each type; level 0 corresponds to no perturbation.
[…]
Gaussian blur TF.gaussian_blur $(r,\sigma,k)=$
$\{(1,0.5,3),(2,1.0,5),$
$(3,1.5,7),(5,2.5,11),$
(10,5.0,21)}(10,5.0,21)\}
JPEG compression PIL.save(..., JPEG quality
format="JPEG", quality) $q\in\{80,50,30,10,1\}$
Random resized crop transforms.RandomResizedCrop scale ranges
$\{(0.85,0.9),(0.7,0.85),$
$(0.5,0.8),(0.3,0.9),$
(0.08,1.0)}(0.08,1.0)\}
Gaussian noise torch.randn_like, torch.clamp $\sigma\in\{0.02,0.05,0.1,0.2,0.5\}$
Perturbation examples can be found in Figs. 17 and 18.
Original Refer to caption Severity level 1 2 3 4 5 Gaussian blur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption JPEG compression Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Random resized crop Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Gaussian noise Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 17: Perturbations example. mandrill.
Original Refer to caption Severity level 1 2 3 4 5 Gaussian blur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption JPEG compression Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Random resized crop Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Gaussian noise Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 18: Perturbations example. Sailboat on lake.
D.6 Additional experimental details
D.6.1 Step size ablation
To assess sensitivity to temporal sampling rate, we vary the frame step size applied to videos already sampled at 8 FPS. For step size $s\in\{1,2,3,4\}$, we subsample every $s$-th frame from all videos, reducing the effective frame rate by a factor of $s$ (yielding 8, 4, 2.67, and 2 FPS respectively). Specifically, this subsampling is applied uniformly to both the VATEX [81] calibration set used to estimate the whitening transform and to the GenVideo [24] test videos being evaluated. We then compute the combined spatial and temporal score using these subsampled sequences. From another perspective, the setup can be viewed as using a larger temporal stride of $s$ between successive differences, without overlap between stride segments in the original video. This ablation tests whether the detector remains robust when operating at lower effective frame rates, which is relevant for computational efficiency. The average AUC results are shown in Fig. 8(a) of the main paper.
D.6.2 FPS ablation
To evaluate robustness to different frame rates, we subsample only the inference videos while keeping the whitening transform fixed (estimated on VATEX [81] at 8 FPS as in all other experiments). We select videos from GenVideo [24] that are originally at 24 FPS with at least 2 seconds duration (425 videos from Gen2 [34], 425 from Lavie [82], and 42 from WildScrape [85], and a matching number of real videos from MSR-VTT [87]), isolating the effect of frame rate from other factors. In contrast to our standard 8 FPS setup, in this experiment we downsample the videos to target frame rates in $\{2,4,8,12,24\}$ FPS using exact subsampling. The downsampling factor and target frame rate are chosen such that
$\frac{\text{current\_fps}}{\text{target\_fps}}=n\in\mathbb{N},$
so the downsampled sequence is obtained by retaining every $n$-th frame. This ensures deterministic frame selection without approximations. The score is computed on the downsampled sequences. Results, summarized in Fig. 8(c) of the main paper, show that performance is essentially unchanged across this range of frame rates, indicating that our method is robust to frame rate variation and that calibrating the whitening transform at 8 FPS does not degrade inference at other frame rates.
D.6.3 Length of video ablation
To evaluate robustness to video length, we follow a procedure similar in nature to the FPS ablation. We select videos from GenVideo [24] that were originally 4 seconds in duration (1380 videos from Gen2 [34], 700 from ModelScope [78], 214 from WildScrape [85], 56 from Sora [62], and 1400 real videos from MSR-VTT [87]), and then truncate each to $\{1,2,3,4\}$ seconds. The whitening transform remains fixed throughout (estimated on VATEX [81] at 2 seconds as in all other experiments). The score is then computed on these truncated videos. Results, shown in Fig. 8(b) of the main paper, demonstrate that our method remains robust across this range of video durations, indicating that calibrating at 2 seconds does not degrade performance on shorter or longer clips.
D.6.4 Backbone encoders ablation
To assess the impact of the feature extractor, we evaluate STALL with five different backbone encoders: three image-based encoders and two video-based encoders. For image encoders, we use DINOv3 [siméoni2025dinov3] (dinov3_vitl16), the lightweight MobileNetV3 [47] (mobilenetv3_large_100 from timm python pacagke), and ResNet-18 [39] (from torchvision.models). For video encoders, we test VideoMAE [74] (MCG-NJU/videomae-base) and ViCLIP [83] (OpenGVLab/ViCLIP-L-14-hf), both from HuggingFace. All encoders are used with pretrained weights, and for image encoders we extract per-frame features and apply them independently to each frame in the video sequence. Results are presented in Table 2 of the main paper.
D.6.5 Calibration set size
To evaluate the sensitivity of our method to the size of the calibration set used for estimating the whitening transform, we systematically vary the number of videos sampled from VATEX [81]. We test calibration set sizes ranging from 1,000 to the full VATEX dataset (33,976 videos) in increments of 1,000 videos, evaluating each configuration across 5 random seeds to ensure statistical reliability. For each calibration set size, we randomly sample the specified number of videos from VATEX, estimate the whitening parameters on this subset, and then evaluate the resulting detector on the GenVideo [24] benchmark. All other pipeline components remain fixed, results are presented in Fig. 7(a) of the main paper.
D.6.6 Calibration set sources
To assess the impact of calibration set composition on detection performance, we evaluate STALL using five different calibration sets drawn from diverse real video sources. We test: (1) VATEX [81] (33,976 videos), (2) GenVideo [40] real subset from MSR-VTT [87] (8,584 videos, corresponding to all MSR-VTT clips excluded from the test set), (3) VideoFeedback [24] combining DiDeMo [3] (1k videos) and Panda70M [25] (1k videos), (4) a combination of Kinetics400 (1,496 videos) and PE [13] (around 1,500 videos from each, see section C.5 for more details), and (5) a balanced hybrid set sampling 1k videos from each of the six sources: MSR-VTT, PE, Panda70M, DiDeMo, VATEX, and Kinetics400. For each calibration set, we estimate the whitening transform using only videos from that set and evaluate the resulting detector on the GenVideo [24] benchmark. All other pipeline components remain fixed. Results are presented in Fig. 6(c) of the main paper.
Table  11 further reports average AUC across all three benchmarks for each calibration source (2K videos each), confirming stable performance across calibration choices.
Benchmark VATEX [81] Kinetics400 [50] DiDeMo [3] Panda-70M [25] MSR-VTT [87] VideoFeedback [40] 0.82 0.82 0.86 0.76 0.73 GenVideo [24] 0.78 0.77 0.76 0.83 0.83 ComGenVid 0.82 0.81 0.87 0.75 0.76
Table 11: Average AUC for different calibration datasets across all three benchmarks.
More broadly, the calibration set defines the reference feature statistics of the detector. If the test domain is not represented in the calibration set (e.g., surveillance or aerial videos), performance may degrade, which is a limitation of the method. Conversely, this also enables domain adaptiveness, as the detector can be adapted to a target domain by choosing an appropriate calibration set. Within the general video regime studied here, we observe stable behavior under calibration and test variation.
D.7 Additional qualitative results
Figure 19: Qualitative examples. Each row shows sampled frames from a video clip, with indicators marking whether its spatial and temporal behavior appears natural or unnatural.
Figure 20: Qualitative examples. Each row shows sampled frames from a video clip, with indicators marking whether its spatial and temporal behavior appears natural or unnatural.
E Efficiency analysis
We conducted a comprehensive efficiency analysis to evaluate the computational performance of all detection methods, measuring model inference time and memory usage under controlled conditions.
E.1 Infrence time analysis
This experiment was performed on a fixed set of 20 videos. Each method ran inference on each video separately (without batching), and we repeated this process 5 times over the same video set to account for performance variability, yielding 100 total inference evaluations per method.
All methods were initialized before timing measurements to ensure a fair comparison. We used Python’s timeit.repeat function to measure execution times, defining the inference time of each method to include video loading. This design ensures that the measured times reflect realistic end-to-end performance, covering both data loading and inference. The complete inference time analysis is provided in Table 12.
Table 12: Inference time comparison for all methods.
Domain Method Mean [sec] Std [sec]
Zero shot images AEROBLADE [68] 2.5266 0.0243
ZED [28] 1.1394 0.0067
RIGID [42] 0.4363 0.0024
Supervised video T2VE [1] 1.9950 0.0102
AIGVdet [5] 5.4216 0.0787
Zero-shot video D3 cos [92] 0.2157 0.0043
D3 L2 [92] 0.2220 0.0015
STALL (ours) 0.2230 0.0010
E.2 memory analysis
To comprehensively evaluate the memory requirements of all methods, we conducted a profiling study that measures both model loading and inference memory consumption. Each method was executed in complete isolation within separate processes to eliminate any potential memory pollution or interference between measurements. We distinguished between two critical phases: (1) model loading memory, which captures the one-time cost of initializing model parameters and loading them onto CPU and GPU, and (2) inference memory, which measures the runtime memory footprint during actual video processing. For each method, we repeated measurements across multiple videos (10 videos × 3 repetitions) to ensure statistical reliability. Memory measurements were captured at two levels: CPU memory was tracked using psutil to monitor RAM consumption, while GPU memory was measured using PyTorch’s [63] CUDA memory tracking facilities to capture peak memory usage. To ensure measurement accuracy, we performed garbage collection and CUDA cache clearing between measurements (gc.collect()) and GPU cache clearing (torch.cuda.empty_cache() and torch.cuda.reset_peak_memory_stats()) performed between each measurement to ensure clean memory states , with deliberate delays to allow the system to stabilize. A detailed analysis of memory consumption appears in Table 13.
Table 13: Memory usage comparison for all methods
domain method model loading cpu (MB) model loading gpu (MB) inference cpu peak (MB) inference gpu peak (MB)
zero shot images AEROBLADE [68] 7875.93 141.02 9711.18 2624.43
ZED [28] 101.11 16.06 1707.72 310.08
RIGID [42] 142.41 327.30 1373.67 567.57
supervised video T2VE [1] 1852.16 1271.43 2795.96 140.96
AIGVdet [5] 488.77 182.59 1849.03 673.18
zero-shot video D3 cos [92] 315.64 1157.72 1470.08 160.30
D3 l2 [92] 315.61 1157.72 1416.19 160.30
STALL (ours) 321.21 1166.77 1647.60 160.30
The memory profiling results reveal significant variations in resource consumption across different detection approaches. Among zero-shot image methods, AEROBLADE [68] demonstrates substantially higher memory footprint during both loading and inference phases, while ZED [28] achieves the most efficient performance. In the supervised video category, T2VE [1] requires notably more GPU memory for model loading compared to AIGVdet [5], though the latter exhibits higher inference-time consumption. Our proposed STALL method, alongside D3 [92], maintains a balanced and efficient memory profile with moderate CPU and GPU usage across both phases, demonstrating comparable efficiency to existing temporal consistency zero-shot method while requiring significantly fewer resources than supervised alternatives. We note that all methods receive video data through CPU memory, which contributes to the observed inference CPU peak measurements across all approaches.